OpenStudy (anonymous):

take derivative of y= square root of (2x+1) by using limit definition? please show me the step. and must use limit definition!

myininaya (myininaya):

\[f(x)=\sqrt{g(x)}\] \[f'(x)=\lim_{h \rightarrow 0}\frac{\sqrt{g(x+h)}-\sqrt{g(x)}}{h} \cdot \frac{\sqrt{g(x+h)}+\sqrt{g(x)}}{\sqrt{g(x+h)}+\sqrt{g(x)}}\] \[=\lim_{h \rightarrow 0} \frac{g(x+h)-g(x)}{(\sqrt{g(x+h)}+\sqrt{g(x)})h}\]

myininaya (myininaya):

then an h will cancel then you will be able to use direct substition

myininaya (myininaya):

\[=\lim_{h \rightarrow 0}\frac{g(x+h)-g(x)}{h} \cdot \lim_{h \rightarrow 0}\frac{1}{\sqrt{g(x+h)}+\sqrt{g(x)}}\]

OpenStudy (anonymous):

i wasnt asking for how the limit definition works just how to do the problem, .... so i can check my answer;

OpenStudy (anonymous):

i know limit definition is f(x+h)-f(x)/h

myininaya (myininaya):

\[=g'(x) \cdot \frac{1}{\sqrt{g(x+0)}+\sqrt{g(x)}}=\frac{g'(x)}{2 \sqrt{g(x)}}\]

OpenStudy (anonymous):

so wehn i plug the functions in, i got 2/h which is not good answer? because it is not integral. so i was asking for what correct steps to do

myininaya (myininaya):

did you not see the first step i did?

myininaya (myininaya):

i multiply topb and bottom by conjugate of top

OpenStudy (anonymous):

yeah it is explaining how the limit definition works. not the problem its just genenrally expalins why the equation created.

OpenStudy (anonymous):

and thats not limit of definition btw. conjugate is just another way to take derivative.

myininaya (myininaya):

the first step i said was to multiply top and bottom by the conjugate of the top then i said an h will cancel and then i said you will be able to use direct substition

OpenStudy (anonymous):

noo, okay this isi what i did and see what i did wrong.

myininaya (myininaya):

your function is f=sqrt(g(x)) where g(x)=2x+!

myininaya (myininaya):

g(x)=2x+1*

OpenStudy (anonymous):

f(x-h)-f(x)/h therefore, square root(2(x-h)+1)-square root(2x+1)/h

OpenStudy (anonymous):

i know how to take the derivative by itself. im just confused because i have to use the limit of definition.

OpenStudy (anonymous):

so i do ^2 for the func. so its 2(x-h)-(2x+1)/h^2 then i mutiply it out, 2x-2h-2x-1/h^2 so all the thing canclees, then i have 2/h... which is wierd?

myininaya (myininaya):

i told you the steps multiply the top and bottom by the conjugate of the top like so: \[\lim_{h \rightarrow 0}\frac{\sqrt{2(x+h)+1}-\sqrt{(2x+1)}}{h} \cdot \frac{\sqrt{2(x+h)+1}+\sqrt{2x+1}}{\sqrt{2(x+h)+1}+\sqrt{2x+1}}\] just like i said above

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{(2(x+h)+1)-(2x+1)}{h(\sqrt{2(x+h)+1}+\sqrt{2x+1})}\]

myininaya (myininaya):

remember i said an h would cancel and then you could use direct substitution

myininaya (myininaya):

\[\lim_{h \rightarrow 0} \frac{2 \not{h}}{\not{h}(\sqrt{2(x+h)+1}+\sqrt{2x+1})}\]

myininaya (myininaya):

now use direct substitution